OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 7 tan^3x - 21 tan x = 0

OpenStudy (loser66):

factor 7tan x out

OpenStudy (anonymous):

ok so i get tan^2x -3

OpenStudy (anonymous):

do i set both to 0

OpenStudy (anonymous):

can you help jdoe?

OpenStudy (jdoe0001):

dunno

OpenStudy (jdoe0001):

well, @Loser66 is correct thus far, you get the common factor out

OpenStudy (jdoe0001):

$$ 7tan^3(x)-21tan(x)=0 \implies 7tan(x)\pmatrix{tan^2(x)-3}=0 $$

OpenStudy (loser66):

to me, 7tan x (tan^2 x -3)=0 ---> tan x =0 or tan^2 x -3 =0 

OpenStudy (loser66):

some more steps to get the answer

OpenStudy (anonymous):

i got 0, pi/3, 2pi/3, pi, 4pi/3, 5pi/3

OpenStudy (jdoe0001):

$$ 7tan^3(x)-21tan(x)=0 \implies 7tan(x)\pmatrix{tan^2(x)-3}=0\\ \\ 7tan(x)=0 \implies tan(x) = 0 \implies tan^{-1}(tan(x)) = tan^{-1}(0)\\ \implies \color{blue}{x = \text{the angle whose tangent is 0 } = \{0,\pi, 2\pi \}}\\ \\ tan^2(x)-3=0 \implies tan^{-1}\pmatrix{tan(x)}=tan^{-1}(\sqrt{3})\\ \implies \color{blue}{x = \text{the angles whose tangent is }\pm\sqrt{3}} $$

OpenStudy (jdoe0001):

so, check your Unit Circle for THOSE angles whose tangent is \(\pm\sqrt{3}\)

OpenStudy (anonymous):

i think i got the right answer.

OpenStudy (anonymous):

see i got another answer without the 0 and pi and it said it was wrong

OpenStudy (jdoe0001):

hmm, yes, is those angles you listed already